• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Fansler

    Log entry time 16:46:32 on June14,2005

    Entry number 146133

    keyword=Techs

    Right Q1 tripped , liquid level read back and 04/16 valve readbacks were zero. Obtained access, saclay rack was off with no errors displayed, power supply indicated no faults, helium level meter displayed 80%. Reset the Saclay rack and ramped up the magnet.
    Liquid level and 04/16 valve readbacks were still zero. Entered the hall again cycled ppower on the helium level meter and discovered an error code. Helium level meter was replaced.

    When we installed the new helium level meter it sent a heat transient to ESR, in turn causing both Dipoles and left Q1 to trip on helium level.

    Helium level retured to left Q1. When trying to ramp left Q1 we appeared to have no control. Recycled power on left Q1, it solved the problem. Helium levels returned in both Dipoles, and ramped normally.

    The ramp problem we noticed on left Q1 was very similar to a problem we had with right Q1 on Saturday. We rebooted the Q1 IOC to solve this problem. We completed the installation of the new Helium level meter on right Q1. Right Q1 is ramped up and operating normally. Both spectrometers are operational.